OpenStudy (anonymous):

solve for w. 5 ( w-3) - 7w = 3

OpenStudy (blurbendy):

5w - 15 - 7w = 3 -2w - 15 =3 -2 w = 18 w = -9

OpenStudy (anonymous):

First, you need to distribute the 5 by multiplying the 5 by everything in the paraenthesis: 5 ( w-3) - 7w = 3 5w-15-7w=3 Now, you need to combine like terms, which will be the 5w and the -7w (Remember: always take the sign that's in the front). -2w-15=3 Now, add 15 to both sides to keep getting that w alone: -2w=18 Now, divide by -2. -2w=18 /-2 /-2 w=-9 Hope this helps. :o)